# Hyperons during proto-neutron star deleptonization and the emission of dark flavored particles

### v1

@@ 5.2 Impact of dark flavored sector on PNS evolution
Figure 6:  Radial profiles of selected quantities, showing the restmass density $\rho$ (top left panels), temperature $T$ (top right panels), electron and muons abundances, $Y_{e}$ and $Y_{\mu}$ (lower left panels), as well as the mass fractions for neutrons ($n$), protons ($p$) and the hyperons $\Lambda$, $\Sigma^{-}$ and $\Xi^{-}$ (lower right panels), at two different post bounce times during the PNS deleptnoization, for the two sets of simulations launched from 18 M (left panels) and 25 M progenitors (right panels), comparing in addition to the NUCLEONS (black lines) and HYPERONS (blue lines) EOS also those with the inclusion of $X^{0}$ losses, denoted as HYPERONS (orange lines).
At around 1 s (top panels of Fig. 6), the abundance of $\Lambda$ reach as high as about $X_{\Lambda}\simeq 1\times 10^{-3}$ at the very center for the 18 M model and about $X_{\Lambda}\simeq 5\times 10^{-3}$ for the 25 M model. Going outwards in radius, the abundances rise, following the temperature profile, reaching their peak at around $X_{\Lambda}\simeq 5\times 10^{-3}$ for the 18 M model and $X_{\Lambda}\simeq 2\times 10^{-2}$ for the 25 M model. Figure 7 shows the corresponding dark photon emissivity (left scale) and luminosity (right scale) profiles. At around 1 s post bounce, dark sector emissivity and luminosity for the 18 M model exceed those of the 25 M. Luminosities reach values of about $10^{52}$ erg for the former mode and about $5\times 10^{51}$ erg for the latter, which is also illustrated in the corresponding evolution of the dark sector luminosity in the top panels of Fig. 5. We attribute this to the smaller branching ratio for the 18 M model, despite having slightly lower $\Lambda$ abundances and temperatures. Note also the sharp drop of the emissivities, after which the luminosities remain constant, which is due to the sudden drop of the $\Lambda$ abundance at around 13.75 km for the 18 M model and at around 14.5 km for the 25 M model, at 1 s post bounce. This is roughly to about one-half of saturation density, corresponding to conditions for the threshold for hyperons to exist within the HYPERON model EOS.
The situation changes during the ongoing PNS deleptonization with the inclusion of the additional losses due to dark photon emission. The $X^{0}$ luminosity evolution is illustrated in Fig. 5 (top panels), from where it becomes clear that the dark-photon losses dominate the cooling at around 1 s post bounce, for both intermediate-mass and high-mass models. However, the excess cooling results in the rapid drop of the PNS temperature, significantly faster than in the reference HYPERONS simulations (see the bottom panels of Fig. 6). As a consequence, the reduced temperature results in lower hyperon abundances, which, in turn cause the dark photon emissivity to drop accordingly. This is illustrated for the 5 s post-bounce curves in Fig. 7 such that the $X^{0}$ luminosities in Fig. 5 decrease below those of the total neutrinos and the later evolution will be determined by the neutrino losses. However, note that this is the case already at around 3 s post bounce for the intermediate-mass 18 M model, while only at around 7 s post bounce for the massive 25 M model. It is related to the generally higher temperatures and central densities for the latter model, featuring generally higher abundances of hyperons.
We note also the feedback from the excess cooling due to $X^{0}$ emission, namely the more rapid drop of the neutrino luminosities and average energies, for all flavors, during the PNS deleptonization phase that is dominated by $X^{0}$ cooling (see Fig. 5). We observe the reduction of the neutrino emission timescale by a factor of roughly two for both intermediate-mass and high-mass models.
@@
Figure 7:  Radial profiles of the dark particle emissivity $\epsilon$ (black solid lines on the left scale) and luminosity $L$ (grey dashed lines on the right scale) at post bounce times of 1 s and 5 s, corresponding to the HYPERONS run, for which the conditions are shown in Fig. 6.
@@ 6 Summary and conclusions
Novel hyperonic EOS are studied in simulations of CCSN in spherical symmetry. These are based on general relativistic neutrino radiation hydrodynamics, featuring six-species Boltzmann neutrino transport. Particular emphasis has been put on the role of hyperons, which have long been studied in the context of cold neutron stars, where it has been speculated that the super-saturation density EOS softens due to the appearance of these additional and heavy degrees of freedom. In the context of CCSN studies, hyperons have so far only been taken into account in failed CCSN explosion studies that result in the formation of black holes [18, 19, 20], however, based on hadronic EOS with a limited set of hyperons. The focus of the present article is on the long-term evolution of neutrino-driven CCSN explosions, i.e. the deleptonization and later cooling phases of the nascent PNS, based on hadronic EOS including a comprehensive set of hyperons that include, $\Lambda$, $\Sigma^{0,\pm}$ and $\Xi^{0,\pm}$. The FSU hyperon EOS employed in the present study is based on the RMF framework [91, 92] with meson-nucleon coupling constants selected to reproduce nuclear matter and finite nuclei properties, to fulfill certain constraints of high dense matter coming from heavy-ion collisions as well as to be consistent with the observation of massive 2 M pulsars.
The results differ from those reported in the comprehensive CCSN hyperon EOS study of Ref. [114] in various ways. The largest difference arises due to the implementation of larger sample of hadronic states in the present paper. These modify the charge neutrality condition such that $Y_{e}+Y_{\mu}\neq Y_{p}$, which is the case in Ref. [114] where no charged hyperons were considered. A second difference is related to the onset densities for the appearance of hyperons, which is the case at somewhat higher densities in Ref. [114], and the abundance of hyperons is significantly higher than for the present HYPERONS EOS. In particular, in Ref. [114] the abundance of $\Lambda$ hyperons exceeds those of the neutrons at high baryon densities on the order of $10^{15}~{}{\rm g}~{}{\rm cm}^{-3}$, even at low temperatures on the order of $1-10~{}{\rm MeV}$.
@@
The implementation of the $\Lambda$ hyperon in the CCSN simulations enables us to study the impact of a possible dark sector. These arise from the flavor-violating transition of strange quarks into down quarks, taken into account here from the decay of $\Lambda$ hyperon into neutron via the emission of dark bosonic degrees of freedom. Other decays, such as that of $\Sigma^{+}$ are neglected due to their generally low abundances under the conditions of high isospin asymmetry encountered at the interior of the PNS.
The excess cooling imposed by the inclusion of dark losses in the CCSN simulations results in a qualitative different evolution, in particular during the long-term PNS deleptonization phase when the abundance of $\Lambda$ hyperons reaches their maximum. It is assumed that these massless particles escape the star once upon production. Possible re-scattering is omitted, which is consistent with the small couplings to the dark sector we are considering [c.f. 73]. As a consequence, the PNS deleptonizes on a shorter timescale due to the different thermal evolution. As a feedback, similar as has been reported for the case of axion cooling (c.f. Ref.s [113, 53, 57]) in the case of light axions from QCD processes, we find that the the neutrino emission timescale is reduced by roughly a factor of two. This is consistent with the argument of Ref. [48], confirming both values of the branching ratios, $10^{-8}$ for the low mass and $10^{-9}$ for the high mass models, as limiting cases corresponding to the intermediate-mass and high-mass PNS models of about 1.6 M and 2.0 M, respectively. We explore these two cases at the level of neutrino-driven CCSN explosions of 18 $M_{\odot}$ and 25 M progenitors from the stellar evolution series of Ref. [111], which result in remnant PNS masses of about 1.6 M and 2.0 M with lower central density and hence lower $\Lambda$ abundance for the former and the opposite for the latter models.
@@
Another simplification applied here is related to the assumption of chemical equilibrium for strangeness imposing equal chemical potentials, e.g., of neutrons and $\Lambda$ hyperons. On the other hand, the CCSN timescales, e.g., weak interactions, mass accretion and diffusion, might be potentially on the same order of magnitude as the timescale for hyperons to appear, however, through weak processes. In other words, chemical equilibrium between strange and non-strange hadrons might not be fulfilled, similar as the condition of weak equilibrium is not fulfilled for muons for simulation times on the order of more than 30 s of the PNS deleptonization phase, as was shown in the Appendix of Ref. [57]. This requires the implementation of non-equilibrium methods in order to solve for the problem of strangeness in CCSN, whihc we leave for future explorations.
@@ Acknowledgments
We would like to thank Pok Man Lo for helpful comments and inspiring discussions. TF acknowledges support from the Polish National Science Centre (NCN) under grant number 2023/49/B/ST9/03941. JMC thanks MICINN for funding through the grant “DarkMaps” PID2022-142142NB-I00. HK acknowledges support from the PRE2020-093558 Doctoral Grant of the spanish MCIN/AEI/10.13039/ 501100011033/. LT acknowledges support from the project CEX2020-001058-M (Unidades de Excelencia “María de Maeztu”) and PID2022-139427NB-I00, financed by the Spanish MCIN/ AEI/10.13039/501100011033/, as well as by the EU STRONG-2020 project, under the program H2020-INFRAIA-2018-1 grant agreement no. 824093, from the DFG through projects No. 315477589 - TRR 211 (Strong-interaction matter under extreme conditions), and from the Generalitat de Catalunya under contract 2021 SGR 171. All scientific computations were performed at the Wroclaw Centre for Scientific Computing and Networking (WCSS).

### v2

@@ 5.2 Impact of dark flavoured sector on PNS evolution
Figure 6:  Radial profiles of selected quantities, showing the restmass density $\rho$ (top left panels), temperature $T$ (top right panels), electron and muons abundances, $Y_{e}$ and $Y_{\mu}$ (lower left panels), as well as the mass fractions for neutrons ($n$), protons ($p$) and the hyperons $\Lambda$, $\Sigma^{-}$ and $\Xi^{-}$ (lower right panels), at two different post bounce times during the PNS deleptnoization, for the two sets of simulations launched from 18 M (left panels) and 25 M progenitors (right panels), comparing in addition to the NUCLEONS (black lines) and HYPERONS (blue lines) EOS also those with the inclusion of $X_{0}$ losses, denoted as HYPERONS (orange lines).
At around 1 s (top panels of Fig. 6), the abundance of $\Lambda$ reach as high as about $X_{\Lambda}\simeq 1\times 10^{-3}$ at the very center for the 18 M model and about $X_{\Lambda}\simeq 5\times 10^{-3}$ for the 25 M model. Going outwards in radius, the abundances rise, following the temperature profile, reaching their peak at around $X_{\Lambda}\simeq 5\times 10^{-3}$ for the 18 M model and $X_{\Lambda}\simeq 2\times 10^{-2}$ for the 25 M model. Figure 7 shows the corresponding dark particle emissivity (left scale) and luminosity (right scale) profiles. At around 1 s post bounce, dark sector emissivity and luminosity for the 18 M model exceed those of the 25 M. Luminosities reach values of about $10^{52}$ erg for the former mode and about $5\times 10^{51}$ erg for the latter, which is also illustrated in the corresponding evolution of the dark sector luminosity in the top panels of Fig. 5. We attribute this to the larger branching ratio for the 18 M model, despite having slightly lower $\Lambda$ abundances and temperatures. Note also the sharp drop of the emissivities, after which the luminosities remain constant, which is due to the sudden drop of the $\Lambda$ abundance at around 13.75 km for the 18 M model and at around 14.5 km for the 25 M model, at 1 s post bounce. This is roughly to about one-half of saturation density, corresponding to conditions for the threshold for hyperons to exist within the HYPERON model EOS.
@@
Figure 7:  Radial profiles of the dark particle emissivity $\epsilon$ (black solid lines on the left scale) and luminosity $L$ (gray dashed lines on the right scale) at post bounce times of 1 s and 5 s, corresponding to the HYPERONS run, for which the conditions are shown in Fig. 6.
The situation changes during the ongoing PNS deleptonization with the inclusion of the additional losses due to dark boson emission. The $X_{0}$ luminosity evolution is illustrated in Fig. 5 (top panels), from where it becomes clear that the dark-boson losses dominate the cooling at around 1 s post bounce, for both intermediate-mass and high-mass models. However, the excess cooling results in the rapid drop of the PNS temperature, significantly faster than in the reference HYPERONS simulations (see the bottom panels of Fig. 6). Now, as the supernova interior features a lower temperature compared to the reference model without $X_{0}$ cooling, the feedback to the hyperon EOS is that the $\Lambda$ abundance is somewhat lower—note the strong temperature dependence of the hyperon EOS discussed in Sec. 3—depending on the timescale during the PNS deleptonization. As a consequence, at some point during the later evolution, the $\Lambda$ abundance will be too low for the process $\Lambda\rightarrow n+X_{0}$ to still operate and the additional cooling due to $X_{0}$ emission ceases, featuring the continuous decrease of the $X_{0}$ luminosity; see the thick dashed lines in the top panels of Figs. 5(a) and 5(b). In a sense, the system self regulates it’s excess cooling, however, depending sensitively on the BR.
This is illustrated for the 5 s post-bounce curves in Fig. 7 such that the $X_{0}$ luminosities in Fig. 5 decrease below those of the total neutrinos and the later evolution will be determined by the neutrino losses. However, note that this is the case already at around 3 s post bounce for the intermediate-mass 18 M model, while only at around 7 s post bounce for the massive 25 M model. It is related to the generally higher temperatures and central densities for the latter model, featuring generally higher abundances of hyperons.
We note also the feedback from the excess cooling due to $X_{0}$ emission, namely the more rapid drop of the neutrino luminosities and average energies, for all flavours, during the PNS deleptonization phase that is dominated by $X_{0}$ cooling (see Fig. 5). We observe the reduction of the neutrino emission timescale by a factor of roughly two for both intermediate-mass and high-mass models.
@@ 6 Summary and conclusions
Novel hyperonic EOS are studied in simulations of CCSN in spherical symmetry. These are based on general relativistic neutrino radiation hydrodynamics, featuring six-species Boltzmann neutrino transport. Particular emphasis has been put on the role of hyperons, which have long been studied in the context of cold neutron stars, where it has been speculated that the super-saturation density EOS softens due to the appearance of these additional and heavy degrees of freedom. In the context of CCSN studies, hyperons have so far only been taken into account in failed CCSN explosion studies that result in the formation of black holes [26, 29, 30], however, based on hadronic EOS with a limited set of hyperons. The focus of the present article is on the long-term evolution of neutrino-driven CCSN explosions, i.e. the deleptonization and later cooling phases of the nascent PNS, based on hadronic EOS including a comprehensive set of hyperons that include, $\Lambda$, $\Sigma^{0,\pm}$ and $\Xi^{0,\pm}$. The FSU2H hyperon EOS employed in the present study is based on the RMF framework [103, 104] with meson-nucleon coupling constants selected to reproduce nuclear matter and finite nuclei properties, to fulfil certain constraints of high dense matter coming from heavy-ion collisions as well as to be consistent with the observation of massive 2 M pulsars.
[…]
@@
The implementation of the $\Lambda$ hyperon in the CCSN simulations enables us to study the impact of a possible dark sector. These arise from the flavour-violating transition of strange quarks into down quarks, taken into account here from the decay of $\Lambda$ hyperon into neutron via the emission of dark bosonic degrees of freedom. Other decays, such as that of $\Sigma^{+}$ are neglected due to their generally low abundances under the conditions of high isospin asymmetry encountered at the interior of the PNS.
The excess cooling imposed by the inclusion of dark losses in the CCSN simulations results in a qualitative different evolution, in particular during the long-term PNS deleptonization phase when the abundance of $\Lambda$ hyperons reaches their maximum. It is assumed that these massless particles escape the star once upon production. Possible re-scattering is omitted, which is consistent with the small couplings to the dark sector we are considering (c.f. ref. [83]). As a consequence, the PNS deleptonizes on a shorter timescale due to the different thermal evolution. As a feedback, similar as has been reported for the case of axion cooling (c.f. Refs. [126, 63, 67]) in the case of light axions from QCD processes, we find that the neutrino emission timescale is reduced by roughly a factor of two. This is consistent with the argument of ref. [58], confirming both values of the branching ratios, $10^{-8}$ for the low mass and $10^{-9}$ for the high mass models, as limiting cases corresponding to the intermediate-mass and high-mass PNS models of about 1.6 M and 2.0 M, respectively. We explore these two cases at the level of neutrino-driven CCSN explosions of 18 $M_{\odot}$ and 25 M progenitors from the stellar evolution series of ref. [124], which result in remnant PNS masses of about 1 .6 M and 2.0 M with lower central density and hence lower $\Lambda$ abundance for the former and the opposite for the latter models. We note here that the recent analysis at the level of the average $\bar{\nu}_{e}$-energy and the total $\bar{\nu}_{e}$-energy emitted, deduced from the KAM-II, IBM, BUST and LSD detections of SN1987A [128, 129], indicates that present standard core-collapse supernova models agree quantitatively with these measurements of the first few seconds of events while some tension remains with the late time events of Kam-II, IBM and BUST. The latter is due to generally short cooling timescales of all present long-term supernova simulations, including the PNS deleptonization and cooling phases. Note that considering the KAM-II data alone, current supernova models seemingly overestimate the detected $\bar{\nu}_{e}$ average energy, including the late-time detection. If these finding will be confirmed, in particular also from future supernova neutrino detection of the next galactic events, this would potentially put certain constraints on the requirement of non-standard additional cooling agents, such as the $X_{0}$ losses explored in this work due to the presence of strange hadronic degrees of freedom.
@@
Another simplification applied here is related to the assumption of chemical equilibrium for strangeness imposing equal chemical potentials, e.g., of neutrons and $\Lambda$ hyperons. On the other hand, the CCSN timescales, e.g., weak interactions, mass accretion and diffusion, might be potentially on the same order of magnitude as the timescale for hyperons to appear, however, through weak processes. In other words, chemical equilibrium between strange and non-strange hadrons might not be fulfilled, similar as the condition of weak equilibrium is not fulfilled for muons for simulation times on the order of more than 30 s of the PNS deleptonization phase, as was shown in the Appendix of ref. [67]. This requires the implementation of non-equilibrium methods in order to solve for the problem of strangeness in CCSN, which we leave for future explorations.
@@ Acknowledgments
We would like to thank Pok Man Lo for helpful comments and inspiring discussions. TF acknowledges support from the Polish National Science Centre (NCN) under grant number 2023/49/B/ST9/03941. JMC thanks MICINN for funding through the grant “DarkMaps” PID2022-142142NB-I00 and from the European Union through the grant “UNDARK” of the Widening participation and spreading excellence programme (project number 101159929). HK acknowledges support from the PRE2020-093558 Doctoral Grant of the spanish MCIN/AEI/10.13039/ 501100011033/. LT acknowledges support from the project CEX2020-001058-M (Unidades de Excelencia “María de Maeztu”) and PID2022-139427NB-I00, financed by the Spanish MCIN/ AEI/10.13039/501100011033/, as well as by the EU STRONG-2020 project, under the program H2020-INFRAIA-2018-1 grant agreement no. 824093, from the DFG through projects No. 315477589 - TRR 211 (Strong-interaction matter under extreme conditions), and from the Generalitat de Catalunya under contract 2021 SGR 171. All scientific computations were performed at the Wroclaw Centre for Scientific Computing and Networking (WCSS) under ’grant328’.